OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, -3) in slope-intercept form. Choose one answer. a. y = 3x - 6 b. y = 3x - 3 c. y = 3x + 3 d. y = 3x + 6

OpenStudy (saifoo.khan):

find the slope!

OpenStudy (saifoo.khan):

slope = 3

OpenStudy (saifoo.khan):

y = 3x - 3

OpenStudy (sayan):

m=delta(y)/delta(x) =(-3-6)/(0-3)=-9/-3=3 y-y1=m(x-x1) y-6=3(x-3) y=3x+3 the correct answer is C.

OpenStudy (saifoo.khan):

b. y = 3x - 3

OpenStudy (saifoo.khan):

ITS B!

OpenStudy (saifoo.khan):

the y-intercpt is -3 not + 3

OpenStudy (zarkon):

trivial...plug zero into all the x's...the only function that return a value of -3 is (b)

OpenStudy (sayan):

m=delta(y)/delta(x) =(-3-6)/(0-3)=-9/-3=3 y-y1=m(x-x1) y-6=3(x-3) y=3x-3 the correct answer is B.

OpenStudy (saifoo.khan):

Yay, im correct! :)

OpenStudy (sayan):

Ye, you were right.

OpenStudy (saifoo.khan):

but still u tried, MEDAL for u! :)